OpenStudy (anonymous):

find the prime factorization of 2300

OpenStudy (anonymous):

Divide by the lowest prime or 2: 2300/2 = 1150 Divide by 2 again: 1150/2 = 575 Divide by 5 since it ends in 5: 575/5 = 115 Divide by 5 again: 115/5 = 23 You are done, since 23 is prime. So the prime factors are: 2*2*5*5*23

OpenStudy (anonymous):

2*2*5*5*23

OpenStudy (anonymous):